40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1233     Entry time: Fri Jan 16 18:25:32 2009
Author: Yoichi, Kakeru, Rana 
Type: Update 
Category: LSC 
Subject: Arms were unstable 
The single arm lock had been unstable for both arms in the past few days.

Symptoms:
When an arm was locked by itself, the transmitted power showed a lot of fluctuations (sharp drops).
The first attachment shows the arm power fluctuations in power spectrum and time series.
References are when the boost filters are off for the arm feedback.
You can see that when the boosts are off, the power fluctuates a lot.
Also it is obvious that X-arm is a lot worse than Y.


Diagnosis:
The second attachment is the comparison of the error signal spectra between boosts on and off.
(PD3_I is the error signal of X-arm, PD4_I is Y arm). References are boost on.
Since the arm power fluctuation was suppressed by the gain increase, it was suspected that the main
reason for the power fluctuation is not alignment fluctuation. Rather, it is length or frequency fluctuation.

Then I took spectra and coherences of PD3_I, PD4_I and MC_F with both arms locked independently.
You can see broadband coherence between PD3_I (Xarm) and MC_F (frequency noise). In contrast the coherence
between PD4_I and MC_F is smaller. This means X-arm is more susceptible to the frequency noise than Y.
What can make a simple Fabry-Perot cavity more susceptible to frequency noise ? An offset ?
So I canceled the X-arm offset at the X-arm filter bank. Bingo ! The arm power fluctuation of X-arm became as small as Y-arm
in the dataviewer.
But what is making this offset ?
After watching the dataviewer screen for a while, the arm power fluctuation became larger again. I had to re-adjust the artificial offset
to minimize the fluctuation. This made me think that the source of the offset must be something to do with alignment.
In this case, clipping of the beam at the PD was very suspicious.
So I checked the centering of the POX and POY PDs. As expected, POX was terribly off-centered.
POY was also not exactly at the center of the plateau of DC output.
After centering those PDs, the large offset in the arm loops went away.
Now the arm powers are stable without artificial offset in the loop filters.
The last attachment shows the comparison of arm power fluctuation before and after the PD centering.
(references are the measurements before the centering).
Attachment 1: TRXY.pdf  85 kB  Uploaded Fri Jan 16 13:34:45 2009  | Hide | Hide all
TRXY.pdf
Attachment 2: ErrorSignals.pdf  111 kB  Uploaded Fri Jan 16 15:23:39 2009  | Hide | Hide all
ErrorSignals.pdf
Attachment 3: coherenceBetweenArms.pdf  99 kB  | Hide | Hide all
coherenceBetweenArms.pdf
Attachment 4: ArmPowersAfterPDwasCentered.pdf  83 kB  | Hide | Hide all
ArmPowersAfterPDwasCentered.pdf
ELOG V3.1.3-